OpenStudy (anonymous):

using determinant method solve the equations 4x-3y+20=0 and 3x+2y-2=0

OpenStudy (anonymous):

according to the CoC im not gonna give you the answer... but of course a link i can post for your help! www.math.klte.hu/~muzsnay/Pdf/0607EN/2_determinant.pdf

OpenStudy (callisto):

Haven't done this type of question for long, still give it a try :P 4x-3y+20=0 => 4x -3y = -20 3x+2y-2=0 => 3x +2y = 2 Determinant: |4 -3 | Δ = |3 2 | = (4x2) - (-3x3) = 17 | -20 -3 | Δx = | 2 2 | = -34 x = (Δx) /Δ = -34 / 17 = -2 | 4 -20 | Δy =| 3 2 | = 68 y= (Δy ) / Δ = 68/17 = 4 Got it?

OpenStudy (dumbcow):

@Eric_Playsinacoolway , i don't think you are interpreting it correctly...giving solutions is ok as long as it provided step-by-step help with explanations giving multiple answers with no guidance is what is frowned upon

OpenStudy (anonymous):

okay @dumbcow thanx a lot!